The heat of combustion of sulphur in excess air to produce sulphur(iv)oxide is -297KJmol^-1. Calculate the heat change if 6.4g of sulphur is completely burned in excess air.(S=32)
1
Expert's answer
2020-08-07T14:17:57-0400

While combustion of 1 mol of sulfur 297kJ/mol is released.

From a given mass we found actual number of moles: n=m/M.

n=6.4/32=0.2 (moles).

∆H=0.2×(-297)=-59.4 (kJ/mol).